|Instructions regarding War Diaries and Intelligence
Summaries are contained in F.S. Regs., Vol. 1.
and the Staff Manual respectively Title pages
will be prepared in manuscript.
|Army Form C.2118|
|Unit: 612 Field Squadron R.E.|
|Month and Year: July 1944||(Erase heading not required).||Commanding Officer: Major C.J.C. Bowen|
|BRETTEVILLE L'ORGEUILLEUSE||1||0900||Lt. Kuipers briefed by C.R.E. for road recce on road BRETTEVILLE - CULLY. Sqn. is to maintain this route.|
|1||1400||Part of 1 & 2 Tps, out on road maintenance.|
|1||1800||Lt. Pready brought in darelict enemy lorry, to be use on road-work.|
|1||2030||2 I.C. returned from Rft's. Sec. and Corps Rft. Comp with 2 O.R.'s posted to wait.
Weather: Hot all day, but showers in late afternoon and evening.
|BRETTEVILLE L'ORGUEILLEUSE||2||0830||Part of each Tp. went ou to work on road maint.|
|2||1000||2 I.C. again visited Second Army Adv. Rft's. Sec.|
|2||1100||Warning order received for possible move forward in afternoon. Sqn. at 10 mins notice from 1330hrs; this later lengthened to 2 hrs. notice after O.C.'s. visit to C.R.E.||See Appx "A" attd.|
|2||2030||Offrs. and N.C.O's discussion on the operations of the past week. Followed by Offrs. Conference.|
|3||1030||O.C. attended Sqn. Ldrs. Conference with C.R.E.|
|3||1230||O.C. held conference for 2 I.C., Tp.Ldrs. & S.S.M.: unit likely to remain in present location and carry on with work on MOON route, re-starting tomorrow.
Weather:- Still warm but intermittent rain, very heavy at times. Considerable mud, especially on roads, motor-cycling very hazardous.
|4||Work proceeding on Moon route. 9 tippers, D4, armoured D7 and roller u/Comd. Job 75 to complete at end of day.
All ranks had hot showers (improvised by Hd. Hyg. Sec.) Great morale raiser.
O.C. & S.S.M. visited site at which 1 Tp. Half-Track was loct by enemy action on June 27th:- Five O.R's being still missing. No trace was found of the Half-Track nor of any graves or bodies of the five.
One of the survivors accompanied the O.C.
|4||1600||3 Congers handed in to [ ] Cops Fd. Pk. Coy. R.E.|
|5||2 Tp. continued work on Moon route, as no tipper lorries were available it was not possible usefully to employ more than one Tp.|
|5||1830||F.F.I. inspected by M.O. Favourable report on cleanliness of unit.
No rain all day. Considerable sunshine.
|6||3 Tp. working on Moon route.
1 Tp. and half H.Q. visited 2nd Army Mines Museum in BAYEUX.
Very hot and sunny all day.
|BRETTEVILLE L'ORGUEILLEUSE||7||1 Tp. working on Moon route, 3 Tp. and second half of H.Q. visited 2nd Army Mines Museum in BAYEUX.|
|7||1000||O.C., Lt. Mackay & Lt. Casswell with 5 OR's attended a demonstration of earth-road building organised by C.E. 30 Corps.
2 I.C. agains visited Adv. RA's. Sec. at Army H.Q. and 105 Corps Reinforcement Camp, but no rfts. are yet available for collection although the unit has received posting orders for 11.
|7||3 Half-tracks received replacing those lost in recent operations.|
|8||1 Tp. working on Moon route, 2 Tp. visited 2nd Army Mines Museum in BAYEUX.
News received from England that one Dvr I.C. of Residue nows on X(ii) list.
Weather showery & cooler.
|BRETTEVILLE L'ORGUEILLEUSE||9||2 Tp. finally completed Squadron task on Moon route improvement.
9 OR's Rfts arrived. 1 OR posted to 722 A.W. Coy R.E.
|10||Training programme opened; including 20 mins P.T. & 30 mons Mines daily. 1 OR posted to 13 Fd.Sqn. R.E. Weather cool & showery.|
|11||3 Tp. clearing mines from area map square 9069; total lifted during day = 4 mines (Tmi 35, 35 steel & 42). 1 Tp. constructing range near ST. GABRIEL 8879. O.C. & Lt. Lockyer on recce near ST AUBIN-SUR-MER 0284: camp to be constructed after site cleared of mines. Schu mines suspected.|
|ST. AUBIN-SUR-MER||12||0500||2 Tp. moved to new location ST AUBIN-SUR-MER 024824 and began clearing mines. Days bag was 29 French light A.Tk, 14 S Mines, 4 Tmi 35, estimated 2 more days required to complete clearance.||See Appx 'B' attd.|
|12||0830||O.C., 2 I.C., & Harbour Party arrived ST AUBIN.|
|12||1100||Advance party of tradesmen arrived to be employed on [ ], tiling and generally repairing buildings.|
|12||1230||3 Tp. arrived.|
|12||1430||H.Q. & 1 Tp. arrived. Sqn now complete at ST AUBIN.|
|13||2 Tp. continued mine-clearing. Total bag at end of day: 100 French light A.Tk., 50 S Mines, 4 TMi 35. Mines are being destroyed at low water.
1 & 3 Tps. worked on repairs to buildings required for use by Corps Rest Camp.
|See Appx C attd.|
|14||Weather fine but clowdy, warm.|
|ST. AUBIN-SUR-MER||14||2 Tp. completed clearing of mines in the REVENTORIUM area and destruction of those lifted.||See Appx 'D' attd.|
|15||Sqn. now has Tno Sec8 108 Pnr. Coy, Two Secs, 67 Pw Coy and 66 men of 9 Bn. Gren, Gds, working on 8 Corps rest camp. [
] consists of:
(a) Lifting mines from specified areas completed
(b) Clearing wreckage from beach.
(c) Constructing camp for 30 Offrs, 30 Sgts & 300 OR's.
(d) Repairing 41 houses & making them habitable; 6 of these houses to be completed by 2359 hrs, 16th July.
(b) & (c) are progressing up to schedule. State of (d) is: Clearing 60%, [ ] 80%. Clazing 20%, Latrines 50% Water Supply & [ ] started. [ ] (except for showing [ ] large [ ]) 50%.
|ST. AUBIN-SUR-MER||15||1830||O.C. attended C.R.E. & O Gp.|
|16||0830||O.C. held O. Gp:- 3 Tp. to be reconstituted as Armd Tp. composed of light Sec (in half-tracks) from each Tp. and to come u/comd. 29 Assault Sqn R.E. operating u/comd. 29 Armd. Bde.
1 Tp. with an ordinary sec. from 3 Tp. instead of its own Lt.Sec., comes u/comd. 159 Inf. Bde. Remainder of Sqn to remain u/comd. C.R.E.
|16||1300||1 Tp. moved off for new location 111747.|
|16||1700||3 Tp. moved off for new location 903768.|
|16||2145||R.E. 11 Armd. Div. O.O. N° 2 (Op. Goodwood) received. Work on Corps Rest Camp proceeded during day, reaching approx 80% completion.||See Appx "E" attd.|
|ST. AUBIN-SUR-MER||17||1200||Work on Corps Rest Camp ceased; Approx. completion percentage as follows:- Cleaning 100%, Roofs 95% Glazing 90%, latrines 100% Water supply to abbs & lats 100%. [ ] (except [ ]) 65%|
|LASSON||17||1430||H.Q. & 2 Tp. moved to LASSON 962743 arriving 1600 hrs, Lt. Kuipers sent to move with 13 F.S. R.E. & [ ].|
|17||1800||Report & Sitrep received from 1 Tp.||See Appx F attd.|
|18||2330||Reports received from 3 Tp.P: 9 OR's wounded, one of whom (Sgt. King) feared now dead. 2 Half-tracks are J-casualties. 3 Tp. location with H.Q. 8 R.B. at 098653. No further details.
H.Q. & 2 Tp. remained at LASSON all day expecting orders to move forward. At 2000 hrs, orders came to remain overnight & come under 10 mins. notice to move from 0900 hrs 19th July.
Work dome by detached Tps during day:-
|LASSON||18||0645||1 Tp. filled in ditch across road at 103714 on "Holly" route for wheels.|
|18||1130||3 Tp. improved crossing across rly. cutting using explosive to ramp down the cutting.|
|19||1100||H.Q. & 2 Tp. left LASSON harbour.|
|RANVILLE||19||1400||Arrived new harbour 105729 near RANVILLE.
Recived report from 1 Tp. one O.R. slightly wounded.
|19||1730||O.C. on recce. found the 2 Half-Track casualties reported by 3 Tp.|
|19||1845||- 2200 2 I.C. recovered [...] Half-Tracks to Sqn. harbour.
Weather:- very hot & dusty.
|20||0030||Report received from 3 Tp. who have been destroying enemy eqpt.||See Appx. G. attd.|
|20||1200||1 Tp. working at 067648 on improving routes.|
|20||1615||H.Q. & 2 Tp. moved to new harbour at 096723 near LONGUEVAL.|
|LONGUEVAL||20||1650||Violent storm broke. Thunder & lightening. Torrential rain lasted for 95 minutes leaving mud & puddles everywhere. More heavy showers followed at intervals.|
|20||2150||Rum issue authorised.|
|21||Extremely heavy showers throughout the day with brief spells in between. All roads locally impassable to O.R's. Within the harbour area a jeep, car light recce, Carrier 3-ton 4x4 & Half-Track have all bean bogged at various times although 24hrs ago the ground was bone dry.|
|21||1600||3 Tp., having reverted to Sqn. commend arrived in harbour with H.Q.|
|22||0600||2 & 3 Tps. went out to do route maintenance on route to be used by the Div. in morning West of R.Orne into Corps Reserve.|
|22||1230||H.Q. began to move west. Almost every vehicle had to be towed or winched through the and onto the road.|
|22||1700||H.Q. arrived in new harbour at THAON 970767|
|22||1950||1 Tp. arrived in harbour with H.Q., again under Sqn. Comd.|
|23||0230||F.E. arrived with urgent orders from C.R.E.:- Br. recce. pty. containing 2 Offrs. to report to C.R.E. at 0430, recce. site at LOUVIGNY for Bailey over R.Orne & report again by 0730. O.C. to attend for orders at 0830. Sqn. at 50 mins. notice to move from 1000hrs.|
|23||0350||Lts. Casswell & Kuipers set off with recce pty.|
|23||0720||2 Tp., O.C., and 3 Tp. arrived in harbour in that order, O.C. just in time to get away for CRE's Orders. 2 & 3 Tps very tired, having worked all day & travelled all night.|
|23||1100||2 Tp. lorry which had turned on its side en route during the night was recovered with the aid of a Canadian breakdown lorry.|
|23||1130||1 Tp. moved off to harbour near the br. site & begin work on bankseats, Bridge site at 011654. Tps on normal organisation again (see entry for 0830 hrs, 16th July). 2 & 3 Tps to move to br. site later; two tps 13 Fd. Sqn. R.E. come under comd for bridging.
For infom. regarding this br. op. see Appx H. attached.
|24||H.Q. constructed shower baths, improvised.
1 Cpl. Rft. arrived
|THAON||25||0230||3 Tp. harbour at LOUVIGNY, near the br. site was bombed by enemy aircraft. 4 O.R's wounded. 9 vehs. & 2 motor-cycles damaged.|
|25||1030||3 Tp. arrived back at THAON.|
|25||1500||2 Tp. arrived back at THAON.|
|25||1900||1 Tp. arrived back at THAON.
2 I.C. visited adv. Rft's Sec. (Second Army H.Q.)
31 R.H.U. and 105 C.R.C. returning with 3 Cpl. Rft's.
1 Cpl. Rft. posted to 31 R.H.U.
4 OR's posted from 13 Fd. Sqn. R.E.
|27||Practice for projected assault bridging operation across R.Guigne. Rehearsal carried out on R.Seulles. at 938811 by 2 & 3 Tps.|
|28||0400||2 Tp. (last party) arrived back in harbour from rehearsal. During the day 80% of unit saw cinema show.|
|THAON||28||2100||O.C. & 2 Tp. moved out for a second rehearsal of br. op. This is for trg. only as the project on R.Guigne is off. 1 OR evacuated sick.|
|28||2120||I.O. arrived with movement order. Div is to move to BALLEROY 686690 where it will be at hand Div. of Second Army. Unit heads Div. colm across S.P. at 2902006, 2 Tp. are to continue thir night's bridging & follow on 29th.|
|29||0145||Sqn. less O.C. & 2 Tp. moved out of harbour.|
|BALLEROY||29||0600||Sqn. less O.C. & 2 Tp. arrived new harbour at BALLEROY 686690.|
|29||1835||1 Tp. moved off to come u/comd. 159 Inf. Bde.|
|29||1915||3 Tp. moved off to come u/comd. 29 Armd Bde.|
|29||2045||2 Tp. arrived in harbour with H.Q.|
|30||1215||O.C. & 2 Tp. moved out to new location 695615, near C.R.E.
Weather at present, and for the past week, has been mainly fine and warm. There have been occasional showers but each day has had its quota of scorching hot sunshine.
|30||1830||H.Q. moved to new location 677620 near SALLEN.|
|SALLEN||30||1945||H.Q. arrived 677620.|
|30||2100||Lt. Lockyer came in & reported 2 Tp's only jobs during the day were rd. maint. on points where tks had made deep cuts in the road. He then left again.|
|30||2230||O.C. arrived in harbour.
Locations:- 1 Tp. not known, 2 Tp. 694615, 3 Tp. 692592 (VILLENEUVE).
|30||1100||2 OR's evacuated sick.|
|30 & 31||Work done; - 1 Tp. Rd. repair & culverting 682595.
Rd maint & sweaping for mines 693572.
2 Tp. Rd maint.
3 Tp. Rds & verges cleared in SEPT VENTS 6957 Diversion made from 694590 to 696580. Minefield swept at 696580: 1 TM: 43 & 32 Pieric Pots (A200) lifted.
|SALLEN||31||0800||Report received from 3 Tp.:- "Enemy aircraft dropped a bomb in Tp. harbour about midnight; 7 OR's slighly wounded, 1 veh. wrecked & 2 vehs. damaged. During day 2 OR's evacuated due to battle accidents."|
|31||1 OR wounded (mine explosion) in 3 Tp. and 1 OR. evacuated sick from 1 Tp.|
|SALLEN||31||1820||H.Q. moved forward to 682572 near SEPT VENTS|
|SEPT VENTS||31||1 Tp. at 665535; 2 Tp. 683550; 692592.|
Helaas geen map gevonden.
Source: 11th Armoured Division.
Disclaimer: This War Diary is based on its original, but typos might be corrected. Locations are calculated, so might not be in the correct place. For historical research, always check the originals.